OpenStudy (anonymous):

simplify 5√45 + 3√20

OpenStudy (anonymous):

\[5 \sqrt{45} + 3 \sqrt{20} =5 \sqrt{3 \times 3 \times 5} + 3 \sqrt{2 \times 2 \times 5} \] \[=5 \times 3 \sqrt{ 5} + 3 \times 2 \sqrt{ 5} =15 \sqrt{ 5} + 6 \sqrt{ 5} =21 \sqrt{ 5}\]